OpenStudy (anonymous):

Which of the following points lies on the line whose equation is y = 2x - 3? (0, 3) (1, -1) (-1, 4)

OpenStudy (anonymous):

@sleepyjess

OpenStudy (sleepyjess):

Substitute in the x value for each coordinate and see if it matches the y value

OpenStudy (anonymous):

???

OpenStudy (sleepyjess):

Take the x coordinate from each option and substitute it in for x in the equation

OpenStudy (anonymous):

How am I supposed to know if it equals y?

OpenStudy (sleepyjess):

by substituting in x

OpenStudy (sleepyjess):

Here, I'll show you option A. y = 2x - 3 coordinate = (0, 3) x = 0 y = 2(0) - 3 y = 0 - 3 y = -3 Is the y coordinate -3? No, so A is not correct.

OpenStudy (sleepyjess):

Now just do that for the other 2 options and you will find the answer.

OpenStudy (anonymous):

Where are you having doubts?

OpenStudy (anonymous):

In each option, you are given x and y value.. Just put them in your equation and find out if Left Side = Right Side or not..

OpenStudy (anonymous):

Like, I try first option and you will try other two: (0,3) ----> It means \(x=0, y=3\): So: \(y = 2x-3 \implies 3 = 2(0) - 3 \implies 3 \ne -3\)

OpenStudy (anonymous):

As 3 is not equal to -3, so first option is not right.. Getting?

OpenStudy (anonymous):

Then encourage her to check for other two options left, @sleepyjess

OpenStudy (sleepyjess):

I did...

OpenStudy (anonymous):

No. I did what @sleepyjess said to do and found the answer. I don't just "guess" in school.

OpenStudy (anonymous):

hey, I want to just see your steps for second option, can you now please show that to us or not?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

No problem. I just don't see the point of having to take the time to show my work on this website when I don't even have to do it for school.

OpenStudy (anonymous):

This attitude is neither good nor it is going to help you. We won't tell you the right answer, but we give a hint that, check for both second and third options, otherwise you will end up doing it wrong.. :) Good day..!!!

OpenStudy (anonymous):

I did check the answers. But I don't feel the need to show my work to you two when I don't even have to for my teachers.

OpenStudy (sleepyjess):

We could have already been done by now...

OpenStudy (anonymous):

Right? Lol

OpenStudy (sleepyjess):

I don't see why you don't want to just show to work for the second option. Maybe we are getting a different answer and want to see why you are getting what you are.

OpenStudy (anonymous):

I know I have the correct answer. It doesn't need to go on any further.

OpenStudy (sleepyjess):

Okay, all we wanted to do was check your answer.

OpenStudy (e.mccormick):

How to properly use Desmos to check an answer like this is to graph all parts. Not just the line, but also the points. https://www.desmos.com/calculator/6ttghsuqjp As you can see, only one of the dots is on the line. However, this does not explain the math behind it so it should only be used as a check after you do the math. Why? If you need to take a state test to graduate or need to use this in college, you won't have the web site to use.